OpenStudy (anonymous):

Is this sequence arithmetic? 1, 3, 9, 27, 81... How will I know if a sequence is arithmetic or not?

Parth (parthkohli):

Is \(\rm 2nd ~term - 1st~ term\) equal to \(\rm 3rd~term - 2nd~term\)? If yes, then it's arithmetic.

OpenStudy (anonymous):

it is not arithmetic sequence A sequence is arithmetic if the difference between its any two consecutive terms is same. in your question it could be arithmetic if 3-1=9-3=27-9=81-27 but as they are not equal so it is not arithmetic

OpenStudy (aravindg):

infact this is a "Geometric Progression " with common ratio 3

OpenStudy (anonymous):

Ok. Thanks guys! :)